OpenStudy (firejay5):

Give 3 examples of inverse functions and explain how it's an inverse function. For example 3x - 7.

OpenStudy (firejay5):

@phi

OpenStudy (phi):

first, you should write the function as f(x) = 3x -7 or y = 3x -7 to find the inverse, rename y to x and x to y: x= 3y-7 now solve for y: x= 3y-7 x+7 = 3 y (add 7 to both sides) y= x/3 + 7/3 (divide both sides by 3 you can write this as \[ f^{-1}(x)= \frac{x+7}{3} \]

OpenStudy (phi):

there are well-known inverses: y= x^2 and y= sqrt(x) also cos(x) and acos(x) and e^x and ln(x)

OpenStudy (firejay5):

@phi

OpenStudy (phi):

This will explain it so you can do this problem http://www.khanacademy.org/math/algebra/algebra-functions/function_inverses/v/introduction-to-function-inverses

OpenStudy (firejay5):

Thank! :D